OpenStudy (anonymous):

sin3x= sqrt2/2, 0<&=x<2pie

OpenStudy (anonymous):

\[\sin 3x=\sin \frac{\pi}{4}\]

OpenStudy (anonymous):

how did u figure this out

OpenStudy (anonymous):

some angles must be memorized

OpenStudy (anonymous):

Cause iti s a trigonometric property of the unit circle :) He is right though +1 to him! :D

OpenStudy (anonymous):

Oh! i knew that forgot about the unit circle!! btw thank you!